If P dollars is invested at 100r percent compounded annually, at the end of 2 years it will grow to A=P(1+r)2
At what interest rate will $100 grow to $144.00 in 2 yrs If A=144 and P =100 , find r

Answer by checkley71(8403)

144=100(1+R)^2
144=100(1+2R+R^2)
144=100+200R+100R^2
NOW WE PUT THESE TERMS IN THE QUADRATIC FORMULA FORMAT
100R^2+200R+100-144=0 NOW WE COMBINE THE NUMERIC TERMS
100R^2+200R-44=0 NOW WE EXTRACT A COMMON FACTOR 4
4(25R^2+50R-11)=0 NOW WE FACTOR THE EQUATION
(5X+11)(5X-1)=0 NOW WE SOLVE ONLY THE FACTOR THAT GIVES A POSITIVE ANSWER.
5X-1=0
X=1/5
X=.2 IS THE RATE OF INTEREST OR 20%.